King Waterbeds has an annual cash dividend policy that raises the dividend each year by 5​%. The most recent​ dividend, Div 0 was $ 0.40 per share. What is the​ stock's price if a.  an investor wants a return of 6​%? b.  an investor wants a return of 9​%? c.  an investor wants a return of 11​%? d.  an investor wants a return of 14​%? e.  an investor wants a return of 17​%?

Find the operating cash flow for the year for Robinson and Sons if it had sales revenue of $79,300,000​ cost of goods sold of $35,500,000​ sales and administrative costs of $6,000,000​ depreciation expense of $ 7,000,000 and a tax rate of 30 %
Answer #1

rate positively ..

Ans 1 as per dividend discount model price today = Expected dividend next year /(required rate - growth rate)
Ans a) Price today = 0.4*105%/(6%-5%)
$                 42.00
ans = $                 42.00
Ans b) Price today = 0.4*105%/(9%-5%)
$                 10.50
ans = $                 10.50
Ans c) Price today = 0.4*105%/(11%-5%)
$                   7.00
ans = $                   7.00
Ans d) Price today = 0.4*105%/(14%-5%)
$                   4.67
ans = $                   4.67
Ans e) Price today = 0.4*105%/(17%-5%)
$                   3.50
ans = $                   3.50
Ans 2
Computation of operating cash flow
i Revenue 79,300,000
ii COGS 35,500,000
iii administration cost 6,000,000
iv Depreciation 7,000,000
v=i-ii-iii-iv Profit before tax 30,800,000
vi=v*30% Tax @ 30% 9240000
vii=v-vi Net income 21,560,000
viii=vii+iv Operating cash flow 28,560,000
ans = $        28,560,000